# Towards Automated Semantic Interpretability in Reinforcement Learning via Vision-Language Models

### v1

@@ Abstract
Semantic Interpretability in Reinforcement Learning (RL) enables transparency, accountability, and safer deployment by making the agent’s decisions understandable and verifiable. Achieving this, however, requires a feature space composed of human-understandable concepts, which traditionally rely on human specification and fail to generalize to unseen environments. In this work, we introduce Semantically Interpretable Reinforcement Learning with Vision-Language Models Empowered Automation (SILVA), an automated framework that leverages pre-trained vision-language models (VLM) for semantic feature extraction and interpretable tree-based models for policy optimization. SILVA first queries a VLM to identify relevant semantic features for an unseen environment, then extracts these features from the environment. Finally, it trains an Interpretable Control Tree via RL, mapping the extracted features to actions in a transparent and interpretable manner. To address the computational inefficiency of extracting features directly with VLMs, we develop a feature extraction pipeline that generates a dataset for training a lightweight convolutional network, which is subsequently used during RL. By leveraging VLMs to automate tree-based RL, SILVA removes the reliance on human annotation previously required by interpretable models while also overcoming the inability of VLMs alone to generate valid robot policies, enabling semantically interpretable reinforcement learning without human-in-the-loop.
I Introduction
@@
To learn an interpretable model on a semantically interpretable feature space, various model classes can be considered, including decision trees and rule-based policies, both of which facilitate explicit reasoning about an agent’s actions. Decision trees have been used to represent policies in a structured, human-readable format [4, 5, 6]. Similarly, rule-based methods, such as logical programs and symbolic reasoning approaches, offer an alternative by encoding policies in a form that can be explicitly followed and verified [7, 8]. Additionally, hierarchical decision-making frameworks have been explored, where high-level policies generate interpretable subgoals while low-level controllers handle execution [9, 10]. However, none of these prior works address high-dimensional image or video inputs; instead, they rely either on human-defined semantic concepts or ground-truth feature vectors provided by the environment.
@@
The emergence of Vision-Language Models (VLMs) has opened new possibilities for overcoming these challenges. These models, known for their capabilities in fine-grained visual understanding, language reasoning, and decision-making  [18, 19, 20], have seen widespread use in various tasks  [19, 21]. In this work, we demonstrate that VLMs can substitute humans in defining and extracting semantic features, serving not merely as a “scarecrow in oz”  [22] or a cost-effective fallback to human labor, but as a scalable and automated solution that matches human performance.
@@
We propose Semantically Interpretable Reinforcement Learning with Vision-Language Models Empowered Automation (SILVA), which is to the best of our knowledge the first automated framework that achieves semantic interpretability in RL and generalizes effectively to unseen environments. By leveraging VLMs to extract features from raw environment inputs and training with Interpretable Control Trees, we manage to provide the rationale behind a RL algorithm in terms of human-understandable concepts. Notably, our framework (i) maintains competitive performance with black-box models while surpassing uninterpretable models built on the same feature space, (ii) supports both image and video inputs, effectively capturing temporal information, (iii) generalizes to unseen environments without human-in-the-loop, and (iv) enables computational efficiency by training a lightweight convolutional network as an alternative to VLM-based feature extraction.
II Related Work
II-A Interpretable Reinforcement Learning
@@
Figure 1: Overview of SILVA, our proposed framework. A. Feature Extraction involves a three-step process where: (1) The VLM identifies key features given domain knowledge and environment observation. (2) The VLM generates code to extract features from a single image frame and assesses whether certain key features require multiple frames. If so, it builds upon the single-frame code to extract these features. (3) Input-output pairs are collected using the extractor code and are used to train an Extractor Convolutional Neural Network (Extractor CNN) as a computationally efficient substitute. B. Tree-based RL Training leverages the Extractor CNN to obtain semantic features from image observations, which are subsequently used as inputs to the policy network (an Interpretable Control Tree) to generate control actions for the RL agent.
@@
Interpretable RL is inherently challenging compared to standard supervised learning, as it involves long-term consequences to actions and latent receptions of data through exploration.  [2] A significant line of research has explored interpretability in RL through generating visual saliency maps using Jacobian-based, perturbation-based, and attribution-based methods [23, 13, 24]. However, these approaches primarily focus on post-hoc explanations and fall out of a stricter definition of Interpretable RL  [3], which requires all components including the input space and the decision-making process to be intelligible. Approaches to achieve intelligible decision-making includes interpretable tree-structures, rule and decision lists, and other rule-based methods [3], among which decision trees are arguably the easiest to verify  [25]. Among state-of-the-art tree methods, Differentiable Decision Trees (DDTs) employ a Decision Tree structure that allows for online learning via stochastic gradient descent [25], while Interpretable Continuous Control Trees (ICCTs) extend this capability to continuous control tasks, ensuring both high performance and interpretability [26]. Similarly, POETREE (Policy Extraction through Decision Trees) [27] introduces dynamically growing probabilistic decision trees that adapt throughout training, allowing the tree complexity to evolve based on the task requirements. Our work addresses the limitation of prior tree-based methods by automating feature definition and extraction, thus enabling generalization to unseen environments.
II-B Feature Extraction for Reinforcement Learning
Common techniques that map high-dimensional inputs into low-dimensional feature spaces for RL includes linear nonparametric methods, such as manifold learning and spectral learning, and nonlinear parametric methods, such as deep learning  [12]. Other work focuses on the extraction of visual features that are not inherently connected with semantic meanings.  [23] employs a self-supervised interpretable network (SSINet) to produce fine-grained attention masks for highlighting task-relevant information.  [13] computed perturbation-based saliency of the image input to deduce regions that had the most effect. However, only a limited number of studies have proposed approaches without human-in-the-loop to construct semantic feature spaces for RL training.  [28] leverages information from Atari instruction manuals to construct the feature space, then uses unsupervised object detection and grounding to acquire the corresponding
objects in the environment. However, they observe that current techniques exhibit a lack of reliability  [28]. Our framework distinguishes from prior work by enabling automatic feature definition and extraction without human-in-the-loop.
II-C VLMs for Reinforcement Learning
@@
VLMs have been applied in various aspects of RL training, including reward design, preference feedback, and representation deriving [29, 30, 20]. One prominent approach involves using VLMs to generate reward functions.  [29] use VLMs as zero-shot reward models to specify tasks via natural language, while  [31] leverages VLMs to produce dense reward functions through code generation. Additionally,  [30, 32] use VLMs to provide preference feedback with which they updated the reward functions for preference-based RL. Another line of work leverages VLMs as feature extractors to obtain valuable information which facilitates training.  [20] extract promptable representations from VLMs and decode output tokens into embeddings, which are used to initialize RL policies. However, those representations are not derived for interpretability purposes and might not often be correct. To the best of our knowledge, our work is the first to utilize VLM for creating a semantic feature space for RL training.
III Methodology
@@
We introduce Semantically Interpretable Reinforcement Learning with Vision-Language Models Empowered Automation (SILVA), a framework designed to automatically learn an interpretable decision-making RL policy from an unseen environment. SILVA leverages the visual comprehension and reasoning capabilities of VLMs automate feature extraction through a three-step pipeline: (1) it employs zero-shot prompting on the environment to identify a set of semantic features that deterministically contribute to the decision-making, (2) it generates feature extraction codes that captures both static and temporal information, and (3) it collects a dataset with VLM-generated extraction codes to train a lightweight convolutional neural network (Extractor CNN) as a computationally efficient substitute. Subsequently, utilizing the Extractor CNN, SILVA performs tree-based RL training with an Interpretable Control Tree, establishing a transparent mapping between the semantic feature space and action outputs to derive a semantically interpretable policy for decision-making.
III-A Feature Extraction
@@
The feature extraction process of SILVA involves a three-step pipeline, as shown in Fig. 1. In Step 1, we query the VLM to generate key features that contribute to the decision in the environment. In Step 2, we query the VLM to generate code that extracts the identified features from the environment, either from a single image or a list of images. In Step 3, we collect a dataset from the environment with the VLM-generated extractor code, with which we train an Extractor CNN that will be used in RL training to enable batch computing and increase program speed. The methodology is formalized below.
@@
Step 1: Querying the VLM for Semantic Features. To extract intelligible features from an unseen environment that could be used to train an effective RL policy, we need to first determine what features are useful to model decision-making. In this step, we provide the VLM with:
@@
  • $\mathcal{I}_{0}$: An example image of the environment.
@@
For example, in the Skiing environment, $\mathcal{F}_{s}$ includes features such as the position of skier, the position of next gate, the distance to next gate, the position of nearest tree, and the distance to nearest tree, while $\mathcal{F}_{m}$ includes the current horizontal speed of the skiier. This example is also included in Fig. 1.
Step 2: Querying the VLM for Feature Extraction Code. In this step, we query the VLM to generate feature extraction code, denoted by $\mathcal{G}$, to retrieve the identified set of features $\mathcal{F}$. To mitigate hallucination, we provide visual segmentations for each image, which will be covered in the discussion section. The visual segmentation $\mathcal{S}$ is computed using a segmentation function, represented as $Seg$, such that $\mathcal{S}=\text{Seg}(\mathcal{I})$. The code generation process involves two stages:
Single Image Extraction: For features in $\mathcal{F}$, the VLM generates code $\mathcal{G}_{s}$ to extract as many features as possible from a single image-segmentation pair $(\mathcal{I},\mathcal{S})$. In practice, as demonstrated in our experiments, this approach effectively recovers all features in $\mathcal{F}_{s}$. To enhance the robustness of the generated code, the VLM is provided with multiple example images during this stage to ensure it has been exposed to all possible scenarios. See Algorithm 1.
Multiple Image Extraction: For features in $\mathcal{F}_{m}$, the VLM generates code $\mathcal{G}_{m}$, which extends $\mathcal{G}_{s}$ to process sequential image-segmentation pairs $\{(I_{t-i},S_{t-i}),\dots,(I_{t},S_{t})\}$. One example of such a pair is provided to the VLM during generation. See Algorithm 2.
@@
Algorithm 2 Generating Extraction Code: Image List
@@
Step 3: Dataset Generation and Training an Extractor CNN. In this step, we execute $\mathcal{G}$ on trajectories collected from the environment to generate the feature dataset. For each trajectory, a series of images $\mathcal{T}=\{I_{1},I_{2},\dots,I_{k}\}$ is generated, with corresponding visual segmentations $\mathcal{S}_{t}=\text{Seg}(I_{t})$ for each $t$. The extractor code $\mathcal{G}$ processes each image-segmentation pair $(I_{t},S_{t})$ independently to extract features $\mathbf{v}_{t}=\{v_{1},v_{2},\dots,v_{p}\}$, where $v_{i}\in\mathbb{R}$.
@@
The dataset, consisting of image-feature pairs $\{(\mathcal{I}_{t},\mathbf{v}_{t})\}$, is then used to train an Extractor CNN as a batch-operable and efficient replacement to $\mathcal{G}_{s}$. The trained model predicts semantic features directly from raw images, enabling significantly faster feature extraction during RL training.
@@
If the features require temporal context (i.e., $\mathcal{F}_{m}\neq\emptyset$), $\mathcal{G}_{m}$, consisting of simple mathematical operations on the outputs of $\mathcal{G}_{s}$, is manually adapted to utilize the Extractor CNN in place of $\mathcal{G}_{s}$. For example, in the Skiing environment, $\mathcal{G}_{m}$ retrieves the static features $\mathcal{F}_{s}$ by directly obtaining the values from $\mathcal{G}_{s}$ applied to the last image and acquires the only temporal feature: skiier horizontal speed through subtracting the skiier positions between the first and last image frames. The manual adaptation preserves the exact same logic while ensuring that the subtraction operation is properly batched, significantly accelerating the training process by several orders of magnitude. Without this adaptation, training would be prohibitively time-consuming.
Discussion. Recognizing that hallucination is a well-documented challenge in generative VLM models [33], we provide segmentation along with raw images in Step 2 and 3, using a pre-defined segmentation function $Seg$ that detects objects and their positions in an image by contour and color analysis (details provided in Appendix VII-B). This strategy aligns with common practices in the field [34, 35]. By incorporating segmentation, we mitigate the risk of hallucination and ensure that the generated features and codes are grounded in accurate visual data.
TABLE I: Prompting statistics for VLM-generated code. “Code failure” indicates code that cannot import or run. “Excluded as constant” refers to code that generates invariant outputs across diverse samples and is therefore excluded. “Majority selection coverage” represents the proportion of code yielding identical outputs in the majority voting process. Step 1 refers to the process of generating $\mathcal{G}_{s}$, while Step 2 refers to the process of generating $\mathcal{G}_{m}$.
Environment Type Code Failure Rate (%) Excluded as Constant (%) Majority Selection Coverage (%)
Cart-pole Step 1 0 0 38
Cart-pole Step 2 0 0 100
Skiing Step 1 28 14 26
Skiing Step 2 0 0 100
Boxing Step 1 20 46 14
Boxing Step 2 - - -
TABLE II: Performance of Extractor CNN in extracting feature vectors across different environments, compared with ground truth from VLM-generated code. Continuous components are evaluated with MSE, normalized to [0,1], and scaled by $10^{-3}$; boolean components are evaluated with accuracy. Dashes indicate vacant entries due to varying feature lengths and types across environments.
Environment Feature Vector Indices (1–11)
1 2 3 4 5 6 7 8 9 10 11
Skiing (MSE) 0.11 0.02 0.16 0.20 1.06 2.13 1.93 0.50 1.81
Cart-pole (MSE) 0.008 0.016
Boxing (MSE) 1.07 2.30 23.25 19.42 6.80 0.98 21.71 20.26 12.12
Boxing (Accuracy) 0.993 0.990
III-B Tree-based RL Training

### v2

@@ Abstract
Semantic interpretability in Reinforcement Learning (RL) enables transparency and verifiability by making the agent’s decisions understandable and verifiable. Achieving this, however, requires a feature space composed of human-understandable concepts, which traditionally rely on human specification and may fail to generalize to unseen environments. We introduce interpretable Tree-based Reinforcement learning via Automated Concept Extraction (iTRACE), an automated framework that leverages pre-trained vision-language models (VLM) for semantic feature extraction and interpretable tree-based models for policy optimization. iTRACE first extracts semantically meaningful features, then maps them to policies via interpretable trees. To address the impracticality of running VLMs in RL loops, we distill their outputs into a lightweight model. By leveraging Vision-Language Models (VLMs) to automate tree-based reinforcement learning, iTRACE eliminates the need for human annotation traditionally required by interpretable models, while also addressing the limitations of VLMs alone, such as their lack of grounding in action spaces and inability to directly optimize policies. iTRACE outperforms MLP baselines that use the same interpretable features and matches the performance of CNN-based policies, producing verifiable, semantically interpretable, and human-aligned behaviors without requiring human annotation.
1 Introduction
@@
To learn an interpretable model on a semantically interpretable feature space, various model classes can be considered, including decision trees and rule-based policies, both of which facilitate explicit reasoning about an agent’s actions. Decision trees have been used to represent policies in a structured, human-readable format [30, 26, 18]. Rule-based methods offer an alternative by encoding policies in a form that can be explicitly followed and verified [7, 12]. However, none of these prior works address high-dimensional image or video inputs; instead, they rely either on human-defined semantic concepts or ground-truth feature vectors provided by the environment.
@@
The emergence of Vision-Language Models (VLMs) has opened new possibilities for overcoming these challenges. These models, known for their capabilities in fine-grained visual understanding, language reasoning, and decision-making  [4, 38, 6], have seen widespread use in various tasks  [38, 39]. In this work, we demonstrate that VLMs can substitute humans in defining and extracting semantic features, serving not merely as a cost-effective fallback to human labor, but as a scalable and automated solution that matches human performance.
@@
We propose interpretable Tree-based Reinforcement learning via Automated Concept Extraction (iTRACE), which is to the best of our knowledge the first automated framework that achieves semantic interpretability in RL and generalizes effectively to unseen environments. By leveraging VLMs to extract features from raw environment inputs and training with Interpretable Control Trees, we manage to provide the rationale behind a RL algorithm in terms of human-understandable concepts. Notably, our framework (i) maintains competitive performance with black-box models while surpassing uninterpretable models built on the same feature space, (ii) supports both image and video inputs, effectively capturing temporal information, (iii) generalizes to unseen environments without human-in-the-loop, and (iv) enables computational efficiency by training a lightweight vision model as an alternative to VLM-based feature extraction.
2 Related work
2.1 Interpretable reinforcement learning
@@
Figure 1: Overview of iTRACE, our proposed framework. A. Feature Extraction involves a three-step process where: (1) The VLM identifies key features given domain knowledge and environment observation. (2) The VLM generates code to extract features from a segmentation of a single image frame and determines whether certain key features require multiple frames. If so, it extends the single-frame extraction code to handle multi-frame inputs. The segmentation is performed using the Segment Anything Model (SAM). (3) An exploration policy is used to generate a dataset of feature–raw image frame pairs by applying SAM and the feature extraction code. To avoid the computational overhead of running SAM at every reinforcement learning step, we train an Extractor Module as an efficient substitute. B. Tree-based RL Training leverages the pretrained Extractor Module to obtain interpretable features from image observations, which are subsequently used as inputs to the policy network (an Interpretable Control Tree) to generate control actions for the RL agent.
@@
Interpretable RL is inherently challenging compared to standard supervised learning, as it needs to account for the long-term consequences of actions, partial observability, and sparse feedback during exploration [22]. Post-hoc methods like Jacobian-based, perturbation-based, and attribution-based saliency maps  [25, 10, 34] offer limited interpretability, as they lack transparency in both inputs and decision-making [9]. Approaches to achieve intelligible decision-making include rule lists and decision trees [9], among which decision trees are arguably the easiest to verify  [27]. Among state-of-the-art tree methods, Differentiable Decision Trees (DDTs) support gradient-based optimization of interpretable models [27], while Interpretable Continuous Control Trees (ICCTs) extend this capability to continuous control tasks, balancing performance with interpretability [19]. Similarly, POETREE (Policy Extraction through Decision Trees) [17] introduces dynamically growing probabilistic decision trees that adapt throughout training, allowing the tree complexity to evolve based on the task requirements. Our work addresses the limitation of prior tree-based methods, which previously required manual feature engineering or were limited to structured inputs
2.2 Feature extraction for reinforcement learning
Nonparametric methods (e.g., manifold and spectral learning) and nonlinear parametric methods (e.g., deep learning) are commonly used in RL for feature extraction but typically do not produce interpretable features [15]. Other work focuses on the extraction of visual features that improve performance but do not guarantee human-interpretable concepts. Yet, few works attempt to construct semantically meaningful feature spaces without human supervision for RL training.  [35] leverages information from Atari instruction manuals to construct the feature space, then uses unsupervised object detection and grounding to acquire the corresponding objects in the environment. However, they report that the extracted features could be noisy and inconsistent  [35]. Our framework differs from prior work by enabling automatic feature definition and extraction without human-in-the-loop.
2.3 VLMs for reinforcement learning
@@
VLMs have been applied in various aspects of RL training, including reward design, preference feedback, and representation extraction [21, 33, 6]. One prominent approach involves using VLMs to generate reward functions.  [21] use VLMs as zero-shot reward models to specify tasks via natural language, while  [32] leverages VLMs to produce dense reward functions through code generation. Additionally,  [33, 31] use VLMs to provide preference feedback with which they updated the reward functions for preference-based RL. Another line of work leverages VLMs as feature extractors to obtain valuable information which facilitates training.  [6] extract promptable representations from VLMs and decode output tokens into embeddings, which are used to initialize RL policies. However, these representations are not guaranteed to be accurate or semantically meaningful. To our knowledge, this is the first approach that uses VLMs to automatically construct semantically interpretable feature spaces for RL.
3 Methodology
@@
We introduce interpretable Tree-based Reinforcement learning via Automated Concept Extraction (iTRACE), a framework designed to automatically learn an interpretable decision-making RL policy from an unseen environment. iTRACE leverages the visual comprehension and reasoning capabilities of VLMs automate feature extraction through a three-step pipeline: (1) it employs zero-shot prompting on the environment to identify a set of semantic features that deterministically contribute to the decision-making, (2) it generates feature extraction codes that captures both static and temporal information, and (3) it collects a dataset with VLM-generated extraction codes to train a lightweight vision model (which we refer to as the Extractor Module) as a computationally efficient substitute. Subsequently, utilizing the Extractor Module to replicate VLM feature extraction, iTRACE performs tree-based RL training with an Interpretable Control Tree, establishing a transparent mapping between the semantic feature space and action outputs to derive a semantically interpretable policy for decision-making.
3.1 Feature extraction
@@
The feature extraction process of iTRACE involves a three-step pipeline, as shown in Fig. 1. In Step 1, we query the VLM to generate key features that contribute to the decision in the environment. In Step 2, we query the VLM to generate code that extracts the identified features, either from a single image or a list of images depending on whether the features demand temporal context. Both raw observations and segmentation masks are provided to the VLM in this process. In Step 3, we collect a dataset from the environment with the VLM-generated extractor code, with which we train an Extractor Module to increase program speed in RL training. The methodology is formalized below.
Step 1: Querying the VLM for Semantic Features.
@@
To extract intelligible features from an unseen environment that could be used to train an effective RL policy, we need to first determine what features are useful to model decision-making. In this step, we provide the VLM with:
@@
  • $\mathcal{I}_{0}$: A diverse list of example image observations from the environment.
@@
For example, in the Skiing environment, $\mathcal{F}_{s}$ includes the positions of the next gate (left pole x, right pole x, and y), the closest tree (x, y) and the skier’s x-position, while $\mathcal{F}_{m}$ includes the horizontal speed of the skier.
We instruct the VLM to provide a "requires_temporal_context" field for each feature, enabling it to automatically distinguish between static and temporal features and assign them to $\mathcal{F}_{s}$ or $\mathcal{F}_{m}$. While $\mathcal{F}_{m}$ may be unnecessary in some environments, it is essential in others. Returning to the Skiing example, the skier must actively tilt left and right to maintain speed. Without knowledge of the skier’s motion direction, it is ambiguous which way the skier should tilt. We allow the VLM to determine whether temporal context is required for effective decision-making in a given environment. A complete list of extracted features is provided in Appendix A.2.2.
Step 2: Querying the VLM for Feature Extraction Code.
In this step, we query the VLM to generate feature extraction code, denoted by $\mathcal{G}$, to retrieve the identified set of features $\mathcal{F}$ from one or multiple image-segmentation pairs $(\mathcal{I},\mathcal{S})$. Recognizing that hallucination is a well-documented challenge in generative VLM models [16] (see Appendix. A.5), we use segmentation to provide accurate spatial information and ground the generated extraction code in accurate visual data. This strategy aligns with common practices in the field [20, 37]. The visual segmentation $\mathcal{S}$ is computed using Segment Anything [13], denoted by $Seg$, such that $\mathcal{S}=\text{Seg}(\mathcal{I})$. We use the ViT-H SAM model with its default parameters and no modifications to ensure the generalizability of our approach.
Example Segmentations Labeling: While segmentation improves spatial grounding, its interpretation by the VLM remains prone to hallucination. To ensure the generated code generalizes across diverse environment frames, the VLM must be exposed to a representative set of example images. However, the raw JSON format of segmentation can lead to confusion between segments across different images. To address this, we first prompt the VLM to label a small set of segmentations associated with example frames, processing one image at a time. An example of this labeling process is shown in Fig. 9(b), where all segment labels are generated by the VLM based on raw segmentation masks.
Following this, depending on whether the feature set requires temporal context, the VLM is instructed to take one or multiple image-segmentation pairs $(\mathcal{I},\mathcal{S})$. The labeled segmentations serve as reference examples, helping the VLM infer the most likely attribute values of relevant objects and formulate logic to filter out noisy segments. The code generation process involves two steps:
Single Image Extraction: For features in $\mathcal{F}_{s}$, the VLM generates code $\mathcal{G}_{s}$ to extract their values from a single image-segmentation pair $(\mathcal{I},\mathcal{S})$. To enhance the robustness of the generated code, the VLM is provided with multiple example images and their labeled segmentation during this stage to ensure it has been exposed to most possible scenarios. See Algorithm 1. Also see Fig. 9(c) for an example output on segmentation masks in the Skiing environment.
Multiple Image Extraction: For features in $\mathcal{F}_{m}$, if any, the VLM generates code $\mathcal{G}_{m}$, which extends $\mathcal{G}_{s}$ to process sequential image-segmentation pairs $\{(I_{t-i},S_{t-i}),\dots,(I_{t},S_{t})\}$. Multiple image-segmentation pairs $(\mathcal{I},\mathcal{S})$ are provided to the VLM during generation. See Algorithm 2.
@@
Algorithm 2 Generating Extraction Code: Multiple Images
@@
(a) Original image observation.
(b) VLM-labeled segmentation.
(c) VLM-extracted features.
Figure 2: An example of how raw image observations are processed in our pipeline. (a) Raw image observations are segmented using Segment Anything. (b) A VLM assigns semantic labels to segments in a small set of example images. (c) A VLM then generates code that extracts target features from segmentation masks, applicable to all images. Step (b) is only performed on the example set to guide code generation.
Step 3: Dataset Generation and Training the Extractor Module.
@@
In this step, we execute $\mathcal{G}$ on trajectories collected from the environment to generate the feature dataset. For each trajectory, a series of images $\mathcal{T}=\{I_{1},I_{2},\dots,I_{k}\}$ is generated, with corresponding visual segmentations $\mathcal{S}_{t}=\text{Seg}(I_{t})$ for each $t$. The extractor code $\mathcal{G}$ processes each image-segmentation pair $(I_{t},S_{t})$ independently to extract features $\mathbf{f}_{t}=\{f_{1},f_{2},\dots,f_{n}\}$.
@@
The dataset, consisting of image-feature pairs $\{(\mathcal{I}_{t},\mathbf{f}_{t})\}$, is then used to train an Extractor Module as a batch-operable and efficient replacement to $\mathcal{G}_{s}$. The trained model predicts semantic features directly from raw images, enabling significantly faster feature extraction during RL training. In our experiments, we use a CNN-based Extractor Module (architecture included in Appendix A.3) due to its low computational overhead in simulated environments. However, the framework is compatible with any vision model, including fine-tuned ViT transformers or other task-specific architectures based on deployment needs.
@@
If some features require temporal context (i.e., $\mathcal{F}_{m}\neq\emptyset$), $\mathcal{G}_{m}$, consisting of simple mathematical operations on the outputs of $\mathcal{G}_{s}$, is manually adapted to utilize the Extractor Module in place of $\mathcal{G}_{s}$. For example, in the Skiing environment, $\mathcal{G}_{m}$ calculates skier horizontal speed by subtracting the positions returned by $\mathcal{G}_{s}$ The manual adaptation preserves the exact same logic while ensuring that the subtraction operation is properly batched, significantly accelerating the training process by several orders of magnitude.
3.2 Tree-based RL training